# Entanglement behavior and localization properties in monitored fermion systems

### v1

@@ V.4 Inverse participation ratio and localization properties
As shown in Fig. 8(a) for the $t$-$V$ model and in Fig. 8(b) for the SYK model, the quantity $\overline{\ln({\rm IPR})}$ behaves always linearly with $\ln(\mathcal{N}_{L})$. We have also analyzed the slope $m$ of this linear dependence versus $\gamma$ [see Fig. 8(c)] and, even for this, the behavior is qualitatively the same for all the three cases. From the one side, $m$ changes smoothly with $\gamma$, independently of the integrability properties and of the behavior of the EE. From the other side, we always get a value $-1<m<0$, meaning that the system is never perfectly delocalized nor perfectly localized.
In summary, the models we tested are never localized (as shown in Ref. [83] for the tight-binding case, as they always display an anomalous delocalization akin to a multifractal behavior [104, 105, 65, 41]. We have thus found that, for these monitored systems, localization and delocalization properties seem to have no relation with the entanglement behavior, although the latter may behave very differently.
@@ VI Fermionic ladder model
Finally, we test our fitting function on a slightly different model, which has been introduced and discussed in Refs. [76, 84]. Namely, we consider a system of two coupled fermionic chains, each of them with $L$ sites, interacting via local hopping terms, as shown in Fig. 9. The quadratic Hamiltonian is given by
@@
where $\hat{c}_{j,\sigma}^{(\dagger)}$ are fermionic annihilation (creation) operators on the $j$th site ($j=1,\ldots,L$) of the $\sigma$th chain ($\sigma=1,2$). The hopping amplitudes within the two chains are $t_{1}$ and $t_{2}$, while $t_{12}$ is the interchain hopping amplitude. Each chain is subject to periodic boundary conditions, $\hat{c}^{(\dagger)}_{L+1,\sigma}\equiv\hat{c}^{(\dagger)}_{1,\sigma}$. Chain 1 is referred to as the System, while chain 2 acts as the Ancilla; the global system is referred to as the Ladder, due to the geometry of the coupling. The noise is modeled via random projective measurements of the particle number, $\hat{n}_{j,\sigma}=\hat{c}^{\dagger}_{j,\sigma}\hat{c}_{j,\sigma}$, with measurement probabilities $p_{1}$ and $p_{2}$ for the System and the Ancilla, respectively.
In contrast to what described in the previous sections, here the Ladder undergoes a stroboscopic (and not continuous) evolution, during which the periodic dynamics consists of alternating unitary evolutions and projective measurements [71, 36]. The global system, prepared in a random product state at half-filling, evolves under $\hat{H}_{\rm lad}$ for a time $\tau_{u}$ and is then subject to instantaneous local projective measurements [76, 84]. The cycle repeats $N_{st}$ times until $\tau_{st}=N_{st}\tau_{u}$, at which a steady state is reached (for details on the protocol see Appendix C). The final state of the Ladder is pure $\rho=\ket{\Psi(\tau_{st})}\bra{\Psi(\tau_{st})}$ |, while the density matrix reduced to the System, $\rho_{1}=\text{Tr}_{2}\rho$, obtained by tracing out the Ancilla degrees of freedoms, is generally represented by a mixed state.
Figure 9: Sketch of the fermionic ladder model in Eq. (30). The blue and red spheres indicate the two chains of fermions representing the System and the Ancilla, respectively. Fermions can hop between neighboring sites within the System ($t_{1}$), the Ancilla ($t_{2}$), and between the System and the Ancilla ($t_{12}$). Wavy lines represent noise acting on the System and the Ancilla. After tracing out the Ancilla and partitioning the System into two parts $A$ and $B$, we study the entanglement between them.
Figure 10: (a), (b), (c), (d): The FLN ¯$\overline{\mathcal{E}}_{L/2}$ versus $L$, for different values of $t_{2}$, $p_{1}$, and $p_{2}$. Circles denote the numerical data, while lines correspond to the fitting function. (e), (f), (g), (h): The fit parameter $b$ vs $p_{2}$, for the values of $p_{1}$ and $t_{2}$ of the corresponding panel above. The other parameters used are $t_{1}=1$ and $t_{2}=\pi/2$.
@@
We look at the trajectory-averaged steady-state negativity, by averaging over $N_{\rm r}$ trajectories and also over the last $m=5$ time steps after $\tau_{st}$, in order to smooth out fluctuations. Similarly to Eq. (12), we have
¯$\overline{\mathcal{E}}_{L/2}=\frac{1}{m}\sum^{m}_{s=1}\,\overline{\mathcal{E}_{L/2}(\tau_{st}+s\tau_{u})}\,.$ (32)
@@
$\mathcal{D}_{ij,\sigma\sigma^{\prime}}(\tau)=\braket{\Psi(\tau)}{\hat{c}^{\dagger}_{i,\sigma}\hat{c}_{j,\sigma^{\prime}}}{\Psi(\tau)}\,,$ (33)
@@
This model was studied in Ref. [76] and more extensively in Ref. [84], with the purpose to investigate measurement induced transitions in the presence of non-Markovian noise. A rich phenomenology was observed: for small values of $t_{2}$, a transition from a logarithmic- to an area-law scaling of the entanglement is induced either by $p_{1}$ or by $p_{2}$. On the other hand, for large values of $t_{2}$, the logarithmic behavior persists and is actually enhanced for strong $p_{2}$, so that the Ancilla protects the entanglement of the system from noise. In particular, the logarithmic scaling is clearly seen at larger system sizes $L\gtrsim 80$, with finite-size corrections at lower $L$. In what follows we fix $t_{1}=1$ and $t_{12}=\pi/2$, in order to maximize the coupling between the chains.
In Fig. 10, we show the data for different values of $p_{1}$, $p_{2}$ and $t_{2}$ and the relative fitting curves obtained with Eq. (1), noticing that the numerical data are well described for all the considered parameters. In the top panels we show the FLN vs $L$ for different $t_{2}$, $p_{1}$ (different panels) and $p_{2}$ (different colors). In the bottom panels we show the relative fitting exponent $b$ versus $p_{2}$. In particular, in panel (a) we study the regime of small $t_{2}$ and large $p_{1}$, where the FLN grows at small $L$ and saturates to an area law. This behavior is well fitted by Eq. (1), as also showed by the values of $b$ which are consistently larger than 1, see panel (e). In panels (b) and (c), in correspondence of large $t_{2}$, we observe a regime where $b$ is significantly smaller than one [panels (f) and (g)], corresponding to a regime where the asymptotic FLN scales logarithmically with the system size. Finally for smaller $t_{2}$ and small $p_{2}$ [panel (d)], we distinguish both an area law at large $p_{2}$ and a logarithmic growth at small $p_{2}$, a behavior recalling the same situation of Fig. 1. However, differently from Fig. 1, in this case the exponent is monotonous with the transition parameter, marking a difference between the two cases. Moreover, $b$ is always close to one [panel (h)], making it difficult to locate the exact value of $p_{2}$ corresponding to the crossover between $b>1$ and $b<1$. Indeed, while the analysis based on the fit with Eq. (1) locates the crossover at $p_{2}\approx 0.5$, a refined analysis proposed in Ref. [84] signals the emergence of the transition from a logarithmic to an area phase at smaller values $p_{2}\approx 0.25$.
@@ VII Conclusions
We proposed a function (1) to describe the behavior of the long-time entanglement in monitored fermionic systems, which interpolates between a linear behavior, at small $L$, and power-law behavior, at large $L$. Up to the sizes one can reach with state-of-the-art numerical techniques ($L\sim 10^{1}\div 10^{3}$), we are able to recover a correspondence between the parameters of the function and some entanglement scaling laws already known in literature (from area-law, to logarithmic, subvolume-law, and eventually volume-law behavior).
@@
We tested our function by fitting, in different integrable and nonintegrable models, the asymptotic entanglement entropy attained by evolving under a quantum-state-diffusion dynamics. In particular, we chose three integrable one-dimensional models (namely, the tight binding chain with onsite dephasing, the Kitaev chain both with onsite dephasing and with long-range dissipators) and on two nonintegrable models (namely, the staggered $t-V$ chain and the SYK model). We also tested our function in a ladder fermionic model, finding that it also predicts the asymptotic scaling of the fermionic logarithmic negativity, suggesting that our result is a good indicator of the entanglement scaling, independently of the monotone considered. In all the above cases, we found a good qualitative agreement with the existing knowledge of the entanglement behavior with the system size. Note that the logarithmic growth with $L$, although not explicitly present in our formula of Eq. (1), can be glimpsed by a power-law fitting behavior with an exponent $b\approx 0.8$.
@@
Given the reliability of Eq. (1) in capturing the entanglement behavior for a variety of different models, we think that this result may contribute to the development of the theory of entanglement transitions in monitored systems. We are aware of already existing conformal field theory descriptions of this phenomenon, consistent with a large-$L$ behavior. We think it is however worth investigating whether it would be possible to formulate a theory that can incorporates the small-size behavior not only as “corrections”. Moreover in some cases, as for the description of the intermediate regime in the Kitaev chain with long-range dissipator, the fitting function (1) is likely to perform better than the usual logarithmic scaling guess.
@@
Characterizing the short-size behavior of the entanglement can also be useful from an experimental point of view. In fact, if one could find a way to extrapolate information on the entanglement scaling by looking at the behavior for small sizes, it would be then easier to access any experimental verification with present-day technologies. [114]
@@ Acknowledgements.
We thank M. Fava, I. V. Gornyi, and A. D. Mirlin for fruitful discussions. A. R. acknowledges computational resources from MUR, PON “Ricerca e Innovazione 2014-2020”, under Grant No. PIR01 00011 - (I.Bi.S.Co.). We acknowledge support from the Italian MIUR through PRIN Project No. 2017E44HRF. This work was supported by PNRR MUR project PE0000023-NQSTI. G.C. is supported by ICSC – Centro Nazionale di Ricerca in High-Performance Computing, Big Data and Quantum Computing under project E63C22001000006.
@@ Appendix A Fit stability
Here we provide some arguments regarding the stability of the fit proposed in Eq. (1). We focus on the stability of the parameter $b$, by fitting the same data of Fig. 2 in a range $[L_{\text{min}},L_{\text{max}}]$, with varying $L_{\text{min}}$ and $L_{\text{max}}$.
@@
Figure 12: The behavior of $\overline{S_{L/2}(t)}$ versus $t$ for (a) the staggered tight-binding chain [Eq. (21a), with $W=1$ and $V=0$], (b) the staggered $t$-$V$ chain [Eq. (21a), with $W=V=1$], and (c) the SYK model [Eq. (22)]. We choose $\gamma=0.04$ and report data for different system sizes (see legend). The time step has been fixed as $\delta t=10^{-2}$ and, in all cases, is expressed in units of $J=1$.
@@ Appendix B Time traces
We can observe that the, for the SYK model, the EE saturates the fastest and displays the smallest fluctuations in time. Although in the presence of measurements ($\gamma=0.04$), the SYK model shows a dramatically fast relaxation to the stationary long-time limit value $\overline{S_{L/2}}$ and is self-averaging. On the other hand, the mere absence of integrability does not qualitatively change the main features of the time traces for $\overline{S_{L/2}(t)}$ [compare (a) with (b), where the only difference is to choose $V=0$ or $V=1$ in Eq. (21a), respectively].
@@ Appendix C Fermionic logarithmic negativity in the ladder model
Following Ref. [84], we first show how the correlation matrix (33), for the fermionic ladder model described in Sec. VI, evolves under the combined action of the unitary dynamics and the measurements. Our protocol is composed of (i) a unitary dynamics generated by the Hamiltonian in Eq. (30), $\hat{U}=e^{-i\hat{H}_{\rm lad}\tau_{u}}$, and (ii) a sequence of measurements of the fermionic number $\hat{n}_{j,\sigma}$ on each site, with probability $p_{\sigma}$ (for $\sigma=1,2$).
The state of the ladder after the unitary evolution is given by |$\ket{\Psi(\tau_{u})}=\hat{U}\ket{\Psi(0)}$. Going in the Fourier-Nambu space we can write
@@
^$\hat{\mathbb{R}}_{mn}=\frac{1}{L}\sum_{k}e^{-ik(m-n)}\hat{U}_{k}\,.$ (38)
For what concerns the impact of measurements, the operators $\hat{n}_{l,\mu}$ and $1-\hat{n}_{l,\mu}$ are orthogonal projectors, thus the probability to measure $n_{l,\mu}=1$ is given by $p_{n_{l,\mu}=1}(\tau)=\bra{\Psi(\tau)}\hat{n}_{l,\mu}\ket{\Psi(\tau)}$, while the probability to measure $n_{l,\mu}=0$ is given by $p_{n_{l,\mu}=0}(\tau)=1-p_{n_{l,\mu}=1}(\tau)$. The effect of the measurements translates into the following update rule for the correlation matrix $\mathcal{D}_{ij,\sigma\sigma^{\prime}}(\tau)$ [84]:

### v2

@@ V.4 Inverse participation ratio and localization properties
As shown in Fig. 8(a) for the $t$-$V$ model and in Fig. 8(b) for the SYK model, the quantity $\overline{\ln({\rm IPR})}$ behaves always linearly with $\ln(\mathcal{N}_{L})$. We have also analyzed the slope $m$ of this linear dependence versus $\gamma$ [see Fig. 8(c)] and, even for this, the behavior is qualitatively the same for all the three cases. From the one side, $m$ changes smoothly with $\gamma$, independently of the integrability properties and of the behavior of the EE. From the other side, we always get a value 1<m<0-1<m<0, meaning that the system is neither perfectly delocalized nor perfectly localized.
In summary, in the considered range of $\gamma$, the models we tested are not localized (as shown in Ref. [88] for the tight-binding case), as they always display an anomalous delocalization akin to a multifractal behavior [110, 111, 67, 43]. We have thus found that, for these monitored systems, localization and delocalization properties seem to have no relation with the entanglement behavior, although the latter may behave very differently.
@@ VI Fermionic ladder model
Finally, we test our fitting function on a slightly different model, which has been introduced and discussed in Refs. [78, 90]. Namely, we consider a noninteracting system of two coupled fermionic chains, each of them with $L$ sites, interacting via local hopping terms, as shown in Fig. 9. The quadratic Hamiltonian is given by
$\hat{H}_{\rm lad}=\sum_{j,\sigma}t_{\sigma}\big(\hat{c}^{\dagger}_{j,\sigma}\hat{c}_{j+1,\sigma}+\text{h.c.}\big)+t_{12}\sum_{j}\big(\hat{c}^{\dagger}_{j,1}\hat{c}_{j,2}+\text{h.c.}\big)\,$, (31)
@@
Figure 9: Sketch of the noninteracting fermionic ladder model in Eq. (31). The blue and red spheres indicate the two chains of fermions representing the System and the Ancilla, respectively. Fermions can hop between neighboring sites within the System ($t_{1}$), the Ancilla ($t_{2}$), and between the System and the Ancilla ($t_{12}$). Wavy lines represent noise acting on the System and the Ancilla. After tracing out the Ancilla and partitioning the System into two parts AA and $B$, we study the entanglement between them.
Figure 10: (a), (b), (c), (d): The FLN $\overline{\mathcal{E}}_{L/2}$ versus $L$, for different values of $t_{2}$, $p_{1}$, and $p_{2}$. Circles denote the numerical data, while lines correspond to the fitting function. (e), (f), (g), (h): The fit parameter $b$ vs $p_{2}$, for the values of $p_{1}$ and $t_{2}$ of the corresponding panel above. (i), (j), (k), (l): The length scale $L_{0}$ given in Eq. (2) vs $p_{2}$, for the values of $p_{1}$ and $t_{2}$ of the corresponding panel above. The other parameters used are $t_{1}=1$ and $t_{2}=\pi/2$.
In contrast to what described in the previous sections, here the Ladder undergoes a stroboscopic (and not continuous) evolution, during which the periodic dynamics consists of alternating unitary evolutions and projective measurements [73, 37]. The global system, prepared in a random product state at half-filling, evolves under $\hat{H}_{\rm lad}$ for a time $\tau_{u}$ and is then subject to instantaneous local projective measurements [78, 90]. The cycle repeats $N_{st}$ times until $\tau_{st}=N_{st}\tau_{u}$, at which a steady state is reached (for details on the protocol see Appendix C). The final state of the Ladder is pure $|\rho=\ket{\Psi(\tau_{st})}\bra{\Psi(\tau_{st})}$, while the density matrix reduced to the System, $\rho_{1}=\text{Tr}_{2}\rho$, obtained by tracing out the Ancilla degrees of freedoms, is generally represented by a mixed state.
@@
We look at the steady-state trajectory-averaged negativity, by averaging over $N_{\rm r}$ trajectory realizations and also over the last $m=5$ time steps after $\tau_{st}$, in order to smooth out fluctuations. Similarly to Eq. (13), we have
$\overline{\mathcal{E}}_{L/2}=\frac{1}{m}\sum^{m}_{s=1}\,\overline{\mathcal{E}_{L/2}(\tau_{st}+s\tau_{u})}\,.$ (33)
@@
$\mathcal{D}_{ij,\sigma\sigma^{\prime}}(\tau)=\braket{\Psi(\tau)|\hat{c}^{\dagger}_{i,\sigma}\hat{c}_{j,\sigma^{\prime}}|\Psi(\tau)}\,$, (34)
@@
This model was studied in Ref. [78] and more extensively in Ref. [90], with the purpose to investigate measurement induced transitions in the presence of nonMarkovian noise. A rich phenomenology was observed: for small values of $t_{2}$, a crossover from a logarithmic- to an area-law scaling of the entanglement is induced either by $p_{1}$ or by $p_{2}$. On the other hand, for large values of $t_{2}$, the logarithmic behavior persists and is actually enhanced for strong $p_{2}$, so that the Ancilla protects the entanglement of the system from noise. In particular, the logarithmic scaling is clearly seen at larger system sizes $L\gtrsim 80$, with finite-size corrections at lower $L$. In what follows we fix $t_{1}=1$ and $t_{12}=\pi/2$, to maximize the coupling between the chains.
In Fig. 10, we show the data for different values of $p_{1}$, $p_{2}$ and $t_{2}$ and the relative fitting curves obtained with Eq. (1), noticing that the numerical data are well described for all the considered parameters. In the top panels we show the FLN vs $L$ for different $t_{2}$, $p_{1}$ (different panels) and $p_{2}$ (different colors). In the bottom panels we show the relative fitting exponent $b$ versus $p_{2}$. In particular, in panel (a) we study the regime of small $t_{2}$ and large $p_{1}$, where the FLN grows at small $L$ and saturates to an area law. This behavior is well fitted by Eq. (1), as also showed by the values of $b$ which are consistently larger than 1, see panel (e). In panels (b) and (c), in correspondence of large $t_{2}$, we observe a regime where $b$ is significantly smaller than one [panels (f) and (g)], corresponding to a regime where the asymptotic FLN scales logarithmically with the system size. Finally for smaller $t_{2}$ and small $p_{2}$ [panel (d)], we distinguish both an area law at large $p_{2}$ and a logarithmic growth at small $p_{2}$, a behavior recalling the plots in Fig. 1(a). However, differently from Fig. 1 where no transition exits and one has always an area law with $b=1$, in this case the exponent is monotonous with the transition parameter, marking a difference between the two cases. Moreover, $b$ is always close to one [panel (h)], making it difficult to locate the exact value of $p_{2}$ corresponding to the crossover between $b>1$ and $b<1$. Indeed, while the analysis based on the fit with Eq. (1) locates the crossover at $p_{2}\approx 0.5$, a refined analysis proposed in Ref. [90] signals the emergence of the transition from a logarithmic to an area phase at smaller values $p_{2}\approx 0.25$. In panels (i)-(l) we plot $L_{0}$, i.e., the length scale defined in Eq. (2) and separating the volume-law from the power-law entanglement behavior, versus $p_{2}$. We always find a value of $L_{0}$ smaller than $L_{\rm max}=256$ (the maximum of the range where we apply the fit), confirming that the fit is reliable also in this case.
@@ VII Conclusions
In summary, we have proposed the function in Eq. (1) to describe the behavior of the steady-state long-time EE in monitored fermionic systems, which interpolates between a linear behavior, at small $L$, and a power-law behavior, at large $L$. Up to the sizes one can reach with state-of-the-art numerical techniques ($10^{1}\lesssim L\lesssim 10^{3}$), we are able to recover a correspondence between the parameters of the function and some entanglement scaling laws already known in literature (from area-law, to logarithmic, subvolume-law, and eventually volume-law behavior).
@@
We have tested our function by fitting, in different integrable and nonintegrable models, the steady-state EE attained by evolving under a quantum-state-diffusion dynamics. More specifically, in the nonintegrable cases, we have fitted the steady-state EE versus the coupling $\gamma$ with the environment using a generalized Lorentzian function, and we have recovered the behavior described by Eq. (1). In particular, we have chosen three integrable one-dimensional models (namely, the tight binding chain with onsite dephasing, the Kitaev chain both with onsite dephasing and with long-range dissipators) and on two nonintegrable models (namely, the staggered tVt-V chain and the SYK model). We have also tested our function in a noninteracting fermionic model on a two-leg ladder, finding that it also provides a good description of the scaling of the long-time fermionic logarithmic negativity, suggesting that our result is a good indicator of the entanglement scaling, independently of the monotone considered. In all the above cases, we have found a good qualitative agreement with the existing knowledge of the entanglement behavior with the system size. Note that the logarithmic growth with $L$, although not explicitly present in our formula of Eq. (1), can be glimpsed by a power-law fitting behavior with an exponent $b\approx 0.8$. On the basis of a purely numerical analysis, one cannot rule out that this might also be due to a finite-size effect asymptotically providing an area law.
@@
Let us stress again that, due to the lack of a proper analytical support, the results presented here should not be intended as suitable for predicting any otherwise unknown entanglement phases. However, given the reliability of Eq. (1) in capturing the entanglement behavior for a variety of different models in a fairly wide range of system sizes, we think that this result may contribute to the development of the theory of entanglement transitions in monitored systems. We are aware of already existing conformal field theory descriptions of this phenomenon, consistent with a large-$L$ behavior. We think it is however worth investigating whether it would be possible to formulate a theory that can incorporates the small-size behavior not only as “corrections”. Moreover in some cases, as for the description of the intermediate regime in the Kitaev chain with long-range dissipator, the fitting function (1) is likely to perform better than the usual logarithmic scaling guess.
@@
Characterizing the short-size behavior of the entanglement can also be useful from an experimental point of view. In fact, if one could find a way to extrapolate information on the entanglement scaling by looking at the behavior for small sizes, it would be then easier to access any experimental verification with present-day technologies [120].
@@ Acknowledgements.
We thank M. Fava, I. V. Gornyi, A. D. Mirlin, and M. Szyniszewsi for fruitful discussions. A. R. acknowledges computational resources from MUR, PON “Ricerca e Innovazione 2014-2020”, under Grant No. PIR01 00011 - (I.Bi.S.Co.). We acknowledge support from the Italian MUR through the following projects: PRIN 2017 No. 2017E44HRF and PNRR MUR project PE0000023-NQSTI. G.C. is supported by ICSC – Centro Nazionale di Ricerca in High-Performance Computing, Big Data and Quantum Computing under project E63C22001000006.
@@ Appendix A Fit stability
Here we provide some arguments regarding the stability of the fit proposed in Eq. (1). In particular, we focus on the stability of the parameter $b$, by fitting the same data of Fig. 2 in a range $L_{\text{min}},L_{\text{max}}]$, with varying $L_{\text{min}}$ and $L_{\text{max}}$.
@@
Figure 12: The behavior of $\overline{S_{L/2}(t)}$ versus $t$ for (a) the staggered tight-binding chain [Eq. (22a), with $W=1$ and $V=0$], (b) the staggered $t$-$V$ chain [Eq. (22a), with $W=V=1$], and (c) the SYK model [Eq. (23)]. We choose $\gamma=0.04$ and report data for different system sizes (see legend). The time step has been fixed as $\delta t=10^{-2}$ and, in all cases, is expressed in units of $J=1$. We take $N_{\rm r}=48$ in panels (a) and (c), and $N_{\rm r}=64$ in panel (b).
@@ Appendix B Time traces
We can observe that, for the SYK model, the EE saturates the fastest and displays the smallest fluctuations in time. Although in the presence of measurements ($\gamma=0.04$), the SYK model shows a dramatically fast relaxation to the stationary long-time limit value $\overline{S_{L/2}}$ and is self-averaging. On the other hand, the mere absence of integrability does not qualitatively change the main features of the time traces for $\overline{S_{L/2}(t)}$ [compare (a) with (b), where the only difference is to choose $V=0$ or $V=1$ in Eq. (22a), respectively]. We recall that the results presented in the main text are obtained by further averaging such curves over the time. This double averaging process, over the trajectories and time, allows us to smoothen fluctuations and to get rid of them (we checked that the presented results are stable by further increasing $N_{r}$).
@@ Appendix C Fermionic logarithmic negativity in the ladder model
Following Ref. [90], we first show how the correlation matrix (34), for the free-fermion model on a two leg ladder described in Sec. VI, evolves under the combined action of the unitary dynamics and the measurements. Our protocol is composed of (i) a unitary dynamics generated by the Hamiltonian in Eq. (31), $\hat{U}=e^{-i\hat{H}_{\rm lad}\tau_{u}}$, and (ii) a sequence of measurements of the fermionic number $\hat{n}_{j,\sigma}$ on each site, with probability $p_{\sigma}$ (for $\sigma=1,2$).
[…]
@@
$\hat{\mathbb{R}}_{mn}=\frac{1}{L}\sum_{k}e^{-ik(m-n)}\hat{U}_{k}\,.$ (39)
[…]